OpenStudy (anonymous):

Given f(x) = 3x – 5, what is f(–4)?

OpenStudy (welshfella):

just replace the x i 3x - 5 and work it out

OpenStudy (welshfella):

* x in

OpenStudy (welshfella):

replace the x by -4

OpenStudy (welshfella):

3(-4) - 5 = ?

OpenStudy (welshfella):

3(-4) is 3 times -4 what is 3 times -4 ?

OpenStudy (welshfella):

If you don't interact then i cant help you much...

OpenStudy (anonymous):

Im sorry!!! It's raining here so my WiFi is being stupid.

OpenStudy (anonymous):

17???

OpenStudy (welshfella):

3 * -4 = -12 -12 -5 = -17

OpenStudy (anonymous):

It is a quest for a 4 years old :)